OpenStudy (anonymous):

Evaluate each expression when X= 2, Y= -4, and Z=-7. X-Z

OpenStudy (anonymous):

2-(-7)=9

OpenStudy (mayankdevnani):

X-Z=2-(-7)=2+7=9

OpenStudy (mayankdevnani):

ok @Buc110